# A Benchmark for Data Management in Microservices

### v2

@@ 3.4.2. Continuous Queries
In this section, we describe three different types of continuous queries covering important concerns in Online Marketplace. To specify the queries, we use the syntax of Materialize (mat, [n.d.]), a streaming database. Due to space constraints, we explain one continuous query and the others can be found in our extended version (Laigner et al., 2023).
Online Query #1: Seller Dashboard
The business scenario encountered in a marketplace requires sellers to have and end-to-end overview of the operation in real time to identify trends, such as the popularity of products, and to support decision making, such as when to increase product prices.
@@
SELECT seller_id, COUNT(DISTINCT order_id) as count_orders, COUNT(product_id) as count_items, SUM(total_amount) as total_amount, SUM(freight_value) as total_freight, SUM(total_items - total_amount) as total_incentive, SUM(total_invoice) as total_invoice, SUM(total_items) as total_items FROM order_entries
@@
Online Query #2: Cart Abandonment. A popular use case arising from a series of customer interactions is cart abandonment (Materialize, [n.d.]). A cart is considered abandoned in two cases: (i) prior to checkout submission, and (ii) upon a failed payment processing, if no customer checkout re-submission is identified.
Query Description: In this query, we want to find the cart checkouts that have either failed via stock reservation or payment attempt and have not been involved in a new checkout attempt within the next 10 minutes after the failure. Upon detecting an abandoned cart, a CartAbandoned event is generated for both Cart and Customer microservices. The query specification can be found in our extended version (Laigner et al., 2023).
Online Query #3: Low Stock Warning
To assist sellers, marketplaces usually monitor products’ stock proactively to notify sellers about low inventory and ultimately refrain customers from experiencing unavailability of products.
Query Description: In this query, we want to find the products (and their respective sellers) that are likely to face unavailability in case no replenishment is provided in the near future.
The search is based in the following criteria: The average number of items requested per week in the last month, independently of the result of the reservation and payment, is higher than the present stock level. The Low Stock Warning continuous query specification can be found in our extended version.
@@ 4. Data Management Criteria
4.1. Functional Decomposition
Microservice applications require functional partitioning, i.e. application functionalities are decomposed into different microservices. There are two issues in achieving functional decomposition: (i) how microservices interact and (ii) how microservices manage states.
@@ 4.1. Functional Decomposition
(ii) State management is divided into two categories:
a. Direct data access: Each microservice can directly access its own data (encapsulation principle).
b. Indirect data access: Data within each microservice can only be accessed externally through predefined interfaces (e.g., seller dashboard) or be notified about state changes through events ( e.g., Cart is notified about product updates). It should be noted that this criterion does not require data from different microservices to be stored in different databases.
@@ 4.3. Caching or Replication
4.4. State Management Constraints
4.4.1. Inter-microservices constraints.
Refers to a constraint that cut across microservices.
Constraint #1: Cross-microservice referential integrity
Stock always references an existing product in Product. We take inspiration from the CAP theorem (Gilbert and Lynch, 2002) to define two levels of consistency:
@@ 4.4. State Management Constraints
Constraint #2: No duplicated checkouts. A customer cart belonging to a customer session must not be checked out more than once. This can be safeguarded by having mutual exclusion in each cart. Another way to ensure this constraint is by assigning a customer session ID to each new customer session. The ID is included in the payload of each checkout request submitted to Order. Upon receiving it, Order ensures the same cart checkout does not lead to duplicate order processing.
4.4.2. Intra-microservice constraints.
These include invariants that can be enforced by relying on only local states. They can be achieved through appropriate isolation levels.
Constraint #3: No overselling of items. In Stock microservice, both available and reserved quantities cannot fall below zero for every item. In addition, the reserved quantity must never be higher than the available quantity.
Constraint #4: Maintenance of customer statistics. The system tracks successful and failed payments and deliveries through increments of numbers. An appropriate isolation level is required to ensure that the increment in Customer is not missed.
Constraint #5: Linearized product updates. Updates on each individual product’s price or version in the Product state must be linearized.
Constraint #6: Concurrent Update Delivery transactions must execute in isolation. Since several events can be generated during a Update Delivery transaction, it is essential that concurrent Update Delivery transactions do not operate on the records of the same shipments and packages. The purpose is to avoid emitting duplicated ShipmentNotification and DeliveryNotification events.
Constraint #7: Stock operations triggered by events must execute in isolation. Processing events in Stock requires either serializable isolation or exclusive locks on the items contained in the event payload.
4.5. Event Processing Constraints
@@ 4.6. Performance and Failure Isolation
The functional decomposition of the application allows microservices to operate independently, therefore, providing benefits from the isolation of resources assigned to each microservice. Microservice implementations of Online Marketplace can achieve isolation in two tiers:
@@ 4.7. Logging
Audit logging is a critical concern in applications today as it allows developers to track application events like user activities (e.g., checking out a cart) and state updates (Datadog, [n.d.]). The log recorded during application execution serves as an audit trail, aiding developers in troubleshooting faults and verifying compliance with prescribed business rules.
Logging records of operations is even more pressing in distributed systems, such as microservices. The substantial exchange of events and the complex interplay of independent components make it challenging to reason about errors and failures involving multiple asynchronous microservices.
In Online Marketplace, there are two key events related to logging historical records of operations: ShipmentNotification and PaymentFailed.
@@ 5. Data and Workload Composition
Furthermore, the number of products and sellers should be large enough not to create many conflicts when running. Considering $prob\_s$ and $prob\_p$ as the probabilities of picking a seller and of picking a given seller’s product, respectively, to have less than $\delta$ conflicts, one should pick distributions that lead to $T$ $prob\_s$ $prob\_p$ $\delta$.
Optionally, benchmark users can populate records in other microservices to simulate preexisting data and introduce a degree of overhead in state accesses during the workload execution. We use the size factor to populate the database. A size factor $S$ leads the order table to be initialized with $S$*1K tuples. In consequence, payment, and shipment follow the same size given their tuples refer to an existing order. For tables containing a foreign key attribute referencing the primary key of the above tables, we use the scale factor defined below to define the number of order item tuples per order and consequently, the number of package tuples per shipment.
Ratio of Transactions. It defines the probability of submitting each business transaction and continuous query requests to the system. The transaction ratio can be tuned to specific goals and business scenarios for the benchmark users, such as query- or transaction-heavy scenarios.
Distributions. Distribution in the workload is centered on sellers and their products. For every operation involving a product (add cart item, product price update, and product delete), one has to pick first a seller and then proceed to pick a corresponding item from the seller’s product keyset. In this way, Uniform and Zipfian distributions can be used interchangeably in two cases:
@@ 6. Benchmark Driver
In this section, we discuss the benchmark driver, the program responsible for managing the life-cycle of experiments and collecting metrics as part of Online Marketplace application execution.
@@ 6.1. Driver Functionalities
To manage the life-cycle of an experiment composed by: (i) data generation, (ii) data ingestion, (iii) data check, (iv) workload submission, (v) collection of results, (vi) report generation, and (vii) cleaning of states, we develop a benchmark driver in .NET (Laigner, 2023).
A user specifies the data population, workload composition, ratio of transactions, and distributions through a configuration file. A user usually starts with generating data. Sellers, customers, products, and stock items are generated using Bogus (Chavez, 2024), a popular library for generating synthetic but realistic data, including email, full name, address, and social security number. For products, Bogus generates matching product names and descriptions, including prices and locations. Generated data can be either stored durably or kept in main memory via DuckDB (duc, 2024). Users can load the data from DuckDB and move on to data ingestion, populating the microservices with initial data.
After data generation and ingestion, the driver controls the maximum number of concurrent transactions running in the target platform by: (i) Initializing a number of threads defined by the configuration file. Each thread simulates a user interacting with the application; (ii) Whenever a transaction result returns, the driver pulls a previously used thread from a thread pool and spawns a new transaction submission. Upon experiment completion, the driver computes the metrics of completed transactions and stores the result in a file.
@@ 6.2. Driver Implementation Challenges
The driver enforces the above criteria by maintaining ”idle” customers (those that have either not started any session yet or have completed a session) in a concurrent queue. A customer is pulled from the queue whenever a new customer session starts. Whenever a session terminates, the customer is enqueued again.
@@
Matching Transaction Requests to Asynchronous Results. It is common that platforms for building microservice applications provide results asynchronously (Section  6.1). In such an asynchronous system, the driver must track each submitted transaction and match it with the corresponding asynchronous transaction result to compute the metrics accurately.
To this end, each transaction request is assigned a timestamp and a unique ID. This ID is later used to match the request to a corresponding transaction result that is eventually received. To avoid threads blocking each other, ID generation is decoupled from transaction submission.
In conclusion, a stateful driver is necessary to provide correct transaction input.
7. Example Implementations
To show how Online Marketplace can be used, we implement five different version of Online Marketplace in two competing platforms, Orleans and Statefun, both of which are designed to develop event-driven services with state management functionalities.
7.1. Orleans
@@
Orleans is a framework for building distributed stateful applications. In Orleans, applications are composed of concurrent virtual actors (Bykov et al., 2010), each encapsulating a private state. Developers are offered APIs to log actor state, which allows actors to recover their state upon crashes or server migrations. In our implementation, we use the default Orleans actors’ memory storage API to store the microservices data.
Table 1. Components Design
Actor/Function Description
Cart Models a customer’s cart state and behavior
Customer Models a customer’s state and behavior
Seller Models a seller’s state and behavior
Product Models a product’s state and behavior
Stock Models a stock item’s state and behavior
Order Models a unit of order processing for a single customer and associated state
Payment Models a unit of payment processing for a single customer
Shipment Models a unit of shipment processing for a disjoint group of customers and related state
@@
Due to the single-threaded actor abstraction, developers are encouraged to decompose application functionalities into distinct actors to avoid a few specific virtual actors becoming the bottleneck. Inspired by the guidelines of Wang et al. (Wang et al., 2019), our design aims to maximize parallelism and minimize transaction latency by assigning Online Marketplace functionalities to different actors, as shown in Table 1. We implement Online Marketplace on Orleans 7.2.1, using both the default non-transactional Orleans API and the transactional API, referred to Orleans Transactions (TX) in the rest of the paper. To allow Orleans actors to be reachable from the driver, we deploy an HTTP server on top of the Orleans silo. The server is responsible for parsing and forwarding incoming requests to the appropriate actors, reporting back the transaction results via an HTTP interface.
7.2. Statefun
@@
Statefun is a platform built on top of Flink for running distributed applications based on the concept of stateful functions. Each function encapsulates its own logical state and reacts to incoming messages asynchronously. Each incoming message is processed sequentially, in a way similar to Orleans actors. As a result, applications can be designed by composing functions through messages. Similarly to Orleans, we use the Statefun memory storage API to store the microservices data.
@@
We implement Online Marketplace on Statefun 3.3, and, given the resemblance of both programming models, we opted to model stateful functions with the same design as Orleans. To match the architectural design found in web services, we deploy Statefun with an HTTP ingress. Once the driver submits the transaction input, the ingress acknowledges the reception of the request and dispatches it to the appropriate function. We also deploy an HTTP egress to allow for the collection of transaction results that are eventually completed. The function that terminates a transaction sends a completion message to the egress. The egress operator stores the result and makes it available to clients. This design follows the Statefun documentation (Statefun, 2023b). To execute Stateful Functions runtime, we follow the recommended deployment mode using docker images and the most performant execution style (co-located functions) (Statefun, 2023a).
7.3. Implementation Challenges
@@
Replicating products. Given that both platforms do not provide indexing for the actor or function states, we cannot efficiently query which carts contain a particular product. Therefore, we opted not to implement this replication feature in the experiments.
Continuous queries. We do not use external stream processing systems in order to benchmark only the target platforms rather than the integration of multiple systems. Therefore, we opted to only implement the query dashboard, but not the continuous queries of cart abandonment and low stock warning. The reason is that the order_entries view can be maintained through conventional in-memory data structures and updated through application events, while the others require more complicated stream processing operators such as windowed aggregates and joins, which are not provided natively by the two platforms. This also indicates an opportunity for these platforms to improve their support of microservice applications.
Messaging delivery guarantees. Orleans provides an at-most-once delivery guarantee by default. Although Orleans can be configured to send retries upon timeout, we opted to capture the timeout exception and report to the driver that the transaction has been completed with an error. The reason is that, by enabling retries, the message may arrive multiple times, potentially introducing errors if the application is not idempotent.
On the other hand, Statefun manages state storage and message delivery in an integrated manner, such that in the presence of a delivery error, Statefun transparently retries the delivery up to a timeout and, upon that, rewinds the application to a previously consistent checkpoint (Ewen, 2020). To make the performance results of the two platforms comparable, we disabled checkpointing in Statefun, and in case of a delivery error, we proceeded in the same way as in Orleans.
@@
7.4. A Composite Solution

### v4

@@ 3.4.2. Continuous Queries
In this section, we describe three different types of continuous queries covering important concerns in Online Marketplace. To specify the queries, we use the syntax of Materialize (mat, [n.d.]), a streaming database. Due to space constraints, we explain one continuous query and the others can be found in our extended version (Anonymous, 2024).
Query #1: Seller Dashboard. The business scenario encountered in a marketplace requires sellers to have an end-to-end overview of the operation in real-time to identify trends, such as the popularity of products, and to support decision-making, such as when to increase product prices.
@@
SELECT seller_id, COUNT(DISTINCT order_id) as count_orders, COUNT(DISTINCT seller_id, product_id) as count_items, SUM(total_amount) as total_amount, SUM(freight_value) as total_freight, SUM(total_items - total_amount) as total_incentive, SUM(total_invoice) as total_invoice, SUM(total_items) as total_items
FROM order_entries
@@
Query #2: Cart Abandonment. A popular use case arising from a series of customer interactions is cart abandonment (Materialize, [n.d.]). A cart is considered abandoned in two cases: (i) prior to checkout submission, and (ii) upon a failed payment processing, if no customer checkout re-submission is identified.
Query Description: In this query, we want to find the cart checkouts that have either failed via stock reservation or payment attempt and have not been involved in a new checkout attempt within the next 10 minutes after the failure. Upon detecting an abandoned cart, a CartAbandoned event is generated for both Cart and Customer microservices. The query specification can be found in our extended version (Anonymous, 2024).
Query #3: Low Stock Warning. To assist sellers, marketplaces usually monitor products’ stock proactively to notify sellers about low inventory and ultimately refrain customers from experiencing unavailability of products.
Query Description: In this query, we want to find the products (and their respective sellers) that are likely to face unavailability in case no replenishment is provided in the near future. The search is based in the following criteria: The average number of items requested per week in the last month, independently of the result of the reservation and payment, is higher than the present stock level. The Low Stock Warning continuous query specification can also be found in our extended version (Anonymous, 2024).
@@ 4. Data Management Criteria
4.1. Communication and Data Access
Microservices are functionally partitioned applications (Pritchett, 2008). Two issues arise with this model:
@@ 4.1. Communication and Data Access
(ii) State management is divided into two categories: (a) Direct data access: Each microservice can directly access its own data (encapsulation principle); and (b) Indirect data access: Data within each microservice can only be accessed externally through predefined interfaces (e.g., seller dashboard) or be notified about state changes through events (e.g., Cart is notified about product updates). It should be noted that this criterion does not require data from different microservices to be stored in different databases.
@@ 4.3. Caching or Replication
4.4. State Safety Properties
4.4.1. Inter-microservices Properties
. Refers to a property that cut across microservices.
S#1: Cross-microservice referential integrity. Stock always references an existing product in Product. We take inspiration from the CAP theorem (Gilbert and Lynch, 2002) to define two levels of consistency:
@@ 4.4.1. Inter-microservices Properties
S#2: No duplicate checkouts. The cart of a customer session must not be checked out more than once. This can be safeguarded by having mutual exclusion in each cart. Another way to ensure this property is by assigning a customer session ID to each new customer session. The ID is included in the payload of each checkout request submitted to Order. Upon receiving it, Order ensures the same cart checkout does not lead to duplicate order processing.
4.4.2. Intra-microservice Properties
These include properties that can be enforced by relying on only local states. They can be achieved through appropriate isolation levels.
S#3: No overselling of items. In Stock microservice, both available and reserved quantities cannot fall below zero for every item. In addition, the reserved quantity must never be higher than the available quantity.
S#4: Maintenance of customer statistics. The system tracks successful and failed payments and deliveries through increments of numbers. An appropriate isolation level is required to ensure that the increment in Customer is not missed.
S#5: Linearized product updates. Updates on each individual product’s price or version in the Product state must be linearized.
S#6: Concurrent Update Delivery transactions must execute in isolation. Since several events can be generated during a Update Delivery transaction, it is essential that concurrent Update Delivery transactions do not operate on the records of the same shipments and packages. The purpose is to avoid emitting duplicated ShipmentNotification and DeliveryNotification events.
S#7: Stock operations triggered by events must execute in isolation. Processing events in Stock requires either serializable isolation or exclusive locks on the items contained in the event payload.
4.5. Event Processing Properties
@@ 4.6. Performance and Failure Isolation
The functional partitioning of the application allows microservices to operate independently, offering benefits from the isolation of resources assigned to each microservice (Pritchett, 2008; Kleppmann et al., 2019). Implementations of Online Marketplace can achieve isolation in two tiers:
@@ 4.7. Logging
Audit logging is a critical concern in applications as it allows developers to track application events like user activities (e.g., checking out a cart) and state updates (Datadog, [n.d.]). These events recorded during application execution serves as an audit trail, aiding developers in troubleshooting faults and verifying compliance with prescribed business rules. This is even more pressing in distributed systems, such as microservices. The substantial exchange of events and the complex interplay of independent components make it challenging to reason about errors and failures involving multiple asynchronous microservices. Thus, data systems must store audit events durably. In Online Marketplace, there are two key events related to logging historical records of operations: ShipmentNotification and PaymentFailed.
@@ 5. Data and Workload Composition
Furthermore, the number of products and sellers should be large enough not to create many conflicts when running. Considering $prob\_s$ and $prob\_p$ as the probabilities of picking a seller and of picking a seller’s product, respectively, to have less than $\delta$ conflicts, one should pick distributions that lead to $T$ $prob\_s$ $prob\_p$ $\delta$.
To aid data population, the dataset can be generated based on a given size factor following an uniform distribution. A size factor $S$ leads to 10K customers ($X$), products and also associated stock items ($Y$). Benchmark users can populate records in other microservices to simulate preexisting data and introduce a degree of overhead in state accesses during the workload execution. A size factor $S$ leads the order table to be initialized with $S$*100K tuples uniformly distributed among customers, having the number of order items randomly picked (1-10). In consequence, payment, and shipment follow the same size given their tuples refer to an existing order.
Workload Distributions. Distribution in the workload is centered on sellers and their products. For every operation involving a product (add cart item, product price update, and product delete), one has to pick first a seller and then proceed to pick a corresponding item from the seller’s product keyset. In this way, Uniform and Zipfian distributions can be used interchangeably in two cases:
@@
Metrics. We collect two metrics in this benchmark. Throughput is the number of transactions processed per second, which includes both business transactions that have been successfully completed and continuous queries that have successfully returned a result. End-to-end latency is measured from the moment a client sends a transaction request until the result is received.
Transaction ratios define the probabilities of clients submitting various types of business transactions and continuous queries to the system. We envision it can be set to reflect some potential scenarios:
A. Order-heavy scenarios reflect e-commerce scenarios and high-sales periods, the workload is dominated by order processing (i.e., checkouts and deliveries), having a small percentage of abandoned carts and payments declined. This workload includes both uniform and non-uniform seller distributions, which models for example, some sellers’ products being much more popular than others, leading to orders being canceled due to unavailability. The parameters are inspired by Olist workload (oli, 2024) and specified as product delete with 2%, price update with 3%, seller dashboard with 5%, customer checkout with 30%, and update delivery with 60%.
B. Update-heavy scenarios reflect frequent updates with low-latency requirements, as found in banking, financial trading, and airline applications. This workload models a substantial degree of product updates with non-uniform seller and product distributions. Varying the zipfian constants allows for controlling the number of accesses to Cart and Stock, similar to some extended versions of Smallbank (Shah and Vaz Salles, 2018; Liu et al., 2022). One can set the zipfian constants ranging from 0.2 to 1.2, and transaction ratio as product delete and price update with 30% each, seller dashboard with 2.5%, customer checkout with 12.5%, and update delivery with 25%.
C. Hybrid transactional-analytical scenarios model a substantial number of users continuously querying their data (i.e., the seller dashboard) while the system must also cope with moderate transactional requests. The goal is to prevent the use of simplified caching mechanisms. Thus, seller distribution is uniform while products may have skewed access. Transaction ratios are inspired by HTAP benchmarks (Cole et al., 2011; Zhang et al., 2024) and defined as seller dashboard with 48%, customer checkout with 16%, update delivery with 32%, product delete and update price with 2% each.
@@ 6.1. Driver Functionalities
To manage the life-cycle of an experiment composed by: (i) data generation, (ii) data ingestion, (iii) data check, (iv) workload submission, (v) collection of results, (vi) report generation, and (vii) cleaning of states, we develop a benchmark driver in .NET (Anonymous, 2024).
A user specifies the data population, workload composition, ratio of transactions, and distributions through a configuration file. A user usually starts with generating data. Sellers, customers, products, and stock items are generated using Bogus (Chavez, 2024), a popular library for generating synthetic but realistic data, including email, full name, address, and social security number. For products, Bogus generates matching product names and descriptions, including prices and locations. Generated data can be either stored durably or kept in main memory via DuckDB (duc, 2024). Users can load the data from DuckDB and move on to data ingestion, populating the microservices with initial data. Further details of the driver software, including workload generation, can be found in the extended version of this paper (Anonymous, 2024).
@@ 6.2. Driver Implementation Challenges
To meet this requirement, a thread-safe queue for idle customers is maintained. Whenever a thread initiates a checkout transaction, it picks a customer from the idle queue to represent. After the checkout request is submitted, the thread pushes the customer to the queue, making it available again for other threads. The time a customer lies in the idle queue is driven by the concurrency level supported by the target platform. The faster checkouts are processed, the less time customers spend in the idle queue.
@@
Matching Transaction Requests to Asynchronous Results. It is common that platforms for building microservice applications provide results asynchronously (Section  6.1). In such an asynchronous system, the driver must track each submitted transaction and match it with the corresponding asynchronous transaction result to compute the metrics accurately. To this end, each transaction request is assigned a timestamp and a unique ID. This ID is later used to match the request to a corresponding transaction result that is eventually received. To avoid threads blocking each other, ID generation is decoupled from transaction submission. In conclusion, a stateful driver is necessary to provide correct transaction input.
7. Case Studies
To demonstrate the applicability of Online Marketplace and show the design implications and effects on different systems, we implement and experiment five versions of Online Marketplace in three competing platforms, Orleans, Statefun, and a custom solution. Orleans and Statefun are designed to develop event-driven services with state management functionalities. The custom solution is based on a combination of multiple systems that reflects an usual architecture seek in practice (Laigner et al., 2021; Kleppmann et al., 2019), which will also provide an insight into the complexity faced by the developers. From the implementations and experimental study, we derive lessons learned (referenced by L#) and design decisions (D# in §$\S$ 7.2.3) to foment new systems’ design and improve both state-of-the-art platforms, highlighting the usefulness of Online Marketplace.
7.1. Implementations
7.1.1. Microsoft Orleans
@@
Orleans is a framework for building distributed stateful applications. In Orleans, applications are composed of concurrent virtual actors (Bykov et al., 2010), each encapsulating a private state, that are allocated transparently across machines (orl, 2023a). Developers are offered APIs to log actor state (orl, 2023b), which allows actors to recover their state upon crashes or server migrations. In our implementation, we use the default Orleans actors’ memory storage API to store the microservices data.
@@
Due to the single-threaded actor abstraction, developers are encouraged to decompose application functionalities into distinct actors to avoid a few specific virtual actors becoming the bottleneck. Inspired by the guidelines of Wang et al. (Wang et al., 2019), our design aims to maximize parallelism and minimize transaction latency by assigning Online Marketplace functionalities to different actors. Due to space constraints, further details about the components’ design is found in our extended version (Anonymous, 2024). We implement Online Marketplace on Orleans 7.2.1, using both the default non-transactional Orleans API and the transactional API, referred to Orleans Transactions (TX) in the rest of the paper. To allow Orleans actors to be reachable from the driver, we deploy an HTTP server on top of the Orleans silo. The server is responsible for parsing and forwarding incoming requests to the appropriate actors, reporting back the transaction results via an HTTP interface.
7.1.2. Flink Statefun
@@
Statefun is a platform built on top of Flink for running distributed applications based on the concept of stateful functions (Statefun, 2023d). Flink processes manage the state, handle incoming requests, and invoke the appropriate functions (Statefun, 2023b). Each function encapsulates its own logical state and reacts to incoming messages asynchronously. Each incoming message is processed sequentially, in a way similar to Orleans actors. As a result, applications can be designed by composing functions through messages. Similarly to Orleans, we use the Statefun memory storage API to store the microservices data.
@@
We implement Online Marketplace on Statefun 3.3, and, given the resemblance of both programming models, we opted to model stateful functions with the same design as Orleans. To match the architectural design found in web services, we deploy Statefun with an HTTP ingress. Once the driver submits the transaction input, the ingress acknowledges the reception of the request and dispatches it to the appropriate function. We also deploy an HTTP egress to allow for the collection of transaction results that are eventually completed. The function that terminates a transaction sends a completion message to the egress. The egress operator stores the result and makes it available to clients. This design follows the Statefun documentation (Statefun, 2023c). To execute Stateful Functions runtime, we follow the recommended deployment mode using docker images and the most performant execution style (co-located functions) (Statefun, 2023a).
7.1.3. Implementation Insights
@@
Data replication. Given that both platforms do not provide indexing for the actor or function states, we cannot efficiently query which carts contain a particular product. Although we could carry out this search iterating over all carts of the keyspace, we found this incurs a heavy cost operation, significantly impacting the benchmark analysis. Therefore, we opted not to implement this replication feature in the platforms.
Continuous queries. We do not use external stream processing systems in order to benchmark only the target platforms rather than the integration of multiple systems. Therefore, we opted to only implement the query dashboard, but not the continuous queries of cart abandonment and low stock warning. The reason is that the order_entries view can be maintained through conventional data structures and updated through application events, while the others require more complicated stream processing operators such as windowed aggregates and joins, which are not provided natively by the two platforms.
L#1. Apart from being complex and error-prone, ad-hoc replication and continuous query implementations incur an overhead on both Orleans and Statefun platforms.
Messaging delivery guarantees. Orleans provides an at-most-once delivery guarantee by default. Although Orleans can be configured to send retries upon timeout, we opted to capture the timeout exception and report to the driver that the transaction has been completed with an error. The reason is that, by enabling retries, the message may arrive multiple times, potentially corrupting data if the application is not idempotent (Pritchett, 2008).
On the other hand, Statefun manages state storage and message delivery in an integrated manner, such that in case of a delivery error, Statefun transparently retries the delivery up to a timeout and, upon that, rewinds the application to a previously consistent checkpoint (Ewen, 2020). To make the performance results of the two platforms comparable, we disabled checkpointing in Statefun, and in case of a delivery error, we proceeded in the same way as in Orleans.
L#2. The lack of exactly-once processing guarantees in Orleans requires developers to make their operations idempotent, adding execution overhead and increasing implementation complexity.
@@
Resource Isolation. Both actors in Orleans and functions in Statefun share computational and data storage resources of the node they are allocated to. To isolate the computational resources of a specific microservice, it is necessary to configure a custom grain placement strategy in Orleans (Orleans, 2024a). In Statefun, functions are allocated to Flink Task Managers transparently (Flink, 2024). To make the results comparable, we use Orleans default placement strategy.
L#3. The lack of explicit interfaces for configuring function allocation in Statefun can jeopardize achieving performance and failure isolation for microservices.
7.1.4. A Composite Solution